# Low-pass sampling in Model Predictive Path Integral Control

### v2

@@ III-E Parameters of LP-MPPI
Our proposed algorithm, LP-MPPI, does not introduce any additional parameters beyond those of a low-pass filter. Therefore, in our case, the parameters of the LP-MPPI are the parameters of the Butterworth filter itself. The Butterworth filter has two parameters: cutoff frequency $f_{c}$ and order $o_{\text{LPF}}$. The cutoff frequency controls the width of the passband, which affects the range of frequencies with the highest power density in the drawn control signal perturbations. It affects the trade-off between exploration and exploitation and the bias at lower frequencies, since a lower $f_{c}$ results in a greater bias at lower frequencies and an increased exploitation of low-frequency signals. In turn, a higher $f_{c}$ reduces the bias on the low frequencies and allows exploration in the wider range of signals. The second parameter of the Butterworth filter is its order $o_{\text{LPF}}$, which controls the slope of the power density spectrum in the stopband – roll-off, i.e., how much the attenuation of the signal components grows with the growth of the frequency. The higher the order, the lesser the exploration outside the passband, and the greater the bias towards low frequencies. To better visualize these dependencies, in Figure 3 we present the impact of the filter cutoff frequency $f_{c}$ and its order $o_{\text{LPF}}$ on the magnitude of the frequency response and on the signals in the time domain.
Last but not least, one of the benefits of the proposed approach is the interpretability of the cutoff frequency $f_{c}$ parameter, as it enables intuitive tuning. One of the natural approaches to choosing this parameter is to set it around the natural cutoff frequency of the system to be controlled with LP-MPPI, as inducing higher frequencies requires a significant effort to overcome the damping characteristics of the system itself. Moreover, often, one can even heuristically estimate the highest expected frequency of the system states necessary to maximize the reward or obtain the desired behavior. Furthermore, in some cases, the interpretability of the cutoff frequency may help regularize the system’s behavior, e.g. preventing from exploiting the carelessly designed reward function.
@@
To sum up, the proposed LP-MPPI approach filters out the high-frequency components from the evaluated control signals within the MPPI loop, significantly reducing the high-frequency components in the obtained controls. This reduction is particularly important for robotic systems, as it decreases their wear and biases the search for high-performance control signals towards a more promising area (consider the plausibility of finding a performant control signal in the upper left and bottom right corners of the bottom of Figure 3). Finally, the proposed approach enables one to effectively control the trade-off between exploration and exploitation and offers an intuitive tuning.
@@ IV-A Gymnasium environments
Despite the performance in terms of reward, an important aspect of the control algorithm is the smoothness of the applied controls. To visualize the differences between the considered methods, we present the applied control trajectories for the Ant-v5 environment with $H=15$ and $N=100$ in Figure 6. For fairness of comparison, each algorithm was used with the optimal set of parameters, w.r.t. cumulative reward, determined with Optuna [25] using 100 trials. One can visually assess that the original MPPI algorithm generates undesirably sharp controls. Other baseline methods exhibit significantly better levels of smoothness, comparable to each other and to the proposed approach. To quantitatively compare the smoothness of the signals generated by these methods, we report their Mean Squared Second Derivative (MSSD) and Mean Savitzky-Golay Filter Deviation (MSGFD) in Table I. The results show that our approach achieves very low MSSD and the smallest MSGFD, which confirms that the control trajectories computed with our approach are characterized by a high level of smoothness.
Figure 6: Sample control signals in the Ant-v5 environment for considered MPPI-based control algorithms.
@@ IV-B Quadruped locomotion
In the previous experiment, we showed that the proposed approach outperforms the baselines in some simplified robotic tasks that consider abstract robots. Instead, in this one, we would like to focus on the full-scale simulated quadrupeds with 12DoF. We consider a locomotion task with trot gait, imposed in the reward function to guide the search of the control signals, on two quadrupeds, i.e., Unitree Go2 and MAB Silver Badger (see Figure 4).
Recently, it was shown that MPPI-based approaches can succeed in such complex control tasks by clever adaptation of the standard deviation of the noise distribution and the use of spline interpolation, as it was done in the Dial-MPC approach [14]. We would like to enhance the Dial-MPC with the proposed low-pass filtering (LP Dial-MPC) to evaluate whether it is applicable to different MPPI-based approaches and can improve their performance. Moreover, we would like to compare our method, in this setting, with the best-performing approach from the previous experiment – ColoredMPPI [13]. Thus, we extend the Dial-MPC with colored noise instead of the default white one.
@@
In this experiment, to highlight the robustness of the proposed approach to the choice of its parameters, we do not perform the search with Optuna but instead report the performance for several intuitive parameter sets, i.e. cutoff frequency $f_{c}\in\{2,3,4\}$ for Unitree and $f_{c}\in\{3,4,5\}$ for MAB robot, and orders $o_{\text{LPF}}\in\{2,3,4\}$. We compared their performance with the two above-mentioned baselines, for which we found the sets of the best integer parameters. The results of this experiment can be found in Figure 7. The proposed low-pass filtering approach implemented into the Dial-MPC framework consistently outperforms the default Dial-MPC, by 24% and 41% for the Unitree Go2 and MAB Silver Badger robots, respectively. We attribute these improvements to the more fine-grained control (higher number of decision variables) of the LP Dial-MPC and its ability to directly shape the frequency spectrum of the sampling distribution. In turn, the spectrum shaping capabilities of the colored noise are relatively limited and bias only the lowest frequencies, which results in significantly worse performance (about two times lower rewards than ours).
@@ IV-C Real-world F1TENTH racing
In all previous experiments, we assumed that the models of the controlled systems are perfectly known and are used by the MPPI to search for the best control trajectories. In turn, in this task, we would like to use an analytical model of the F1TENTH car (dynamic single-track model [26] with MF6.1 tire model [27]) and evaluate it under the real-world racing conditions (see Figure 4). The goal of this task is to cover the highest possible distance around the track centerline, on the $14.2\text{\,}\mathrm{m}$ long $1\text{\,}\mathrm{m}$ wide oval racetrack. We defined the cost function by
@@
where $v_{f}$ is the velocity along the centerline, $n$ is the distance to the centerline, $T_{w}$ is the track width, $\alpha$ is slip angle, $\theta$ is the vehicle orientation, and $T_{\theta}$ is the orientation of the centerline. We set the horizon $H=30$, $dt=50$ ms, control frequency to $30\text{\,}\mathrm{Hz}$, and evaluated the algorithms for both the $N=10$ and $N=50$ rollouts. The parameters of all methods were chosen in simulation using Optuna [25] with 50 trials.
@@
Figure 8: Distances covered by the F1TENTH car during $30\text{\,}\mathrm{s}$ long time-trials, controlled by the variants of MPPI algorithms. The proposed LP-MPPI achieves the highest median distances both for $N=10$ and $N=50$.
@@ IV-D Computational overhead
An important aspect of every control method is its computational efficiency. Therefore, we evaluate how much computational overhead the proposed method introduces relative to the nominal MPPI and how it relates to the other considered baselines. To do so, we run all methods for 10 episodes in Ant-v5 (MuJoCo model, $H=15,N=100$) and F1TENTH (compiled analytical model, $H=30,N=10$) environments, using a single core of the Intel Core i5-12500H CPU. We compute the median of the control computation time and relate it to the one obtained by the MPPI. In Table II, we present the results of this experiment. One can see that all baselines introduce some notable computational overhead in the F1TENTH environment, since the compiled analytical model, which is responsible for most of the computations, is very fast. Note that the proposed method introduces the second smallest overhead of 2.4%. In turn, in the case of a relatively heavy dynamics model, e.g., Ant-v5 environment, we observe some counterintuitive results, like the decrease in the compute time for LP-MPPI and ColoredMPPI. We suppose that this may be caused by the variability in the timings, due to the use of a standard OS instead of the real-time one, or be an effect of filtering out the higher frequencies from the control signal, which may simplify the underlying physics simulation. In summary, a computationally intense dynamics evaluation causes the overhead introduced by the proposed method to be negligible.
@@ V Conclusions
In this work, we introduced Low-Pass Model Predictive Path Integral Control (LP-MPPI), a novel, easy-to-implement enhancement to MPPI that incorporates low-pass filtering into the sampling process. By directly shaping the frequency spectrum of control trajectory perturbations, LP-MPPI eliminates harmful high-frequency noise and improves the efficiency of the search for optimal control trajectories. Unlike existing smoothing techniques or colored noise sampling, our approach offers fine-grained control over the balance between exploration and exploitation in the frequency domain, making it highly adaptable to various robotic systems.

### v3

@@ III-E Parameters of LP-MPPI
Our proposed algorithm, LP-MPPI, does not introduce any additional parameters beyond those of a low-pass filter. Therefore, in our case, the parameters of the LP-MPPI are the parameters of the Butterworth filter itself – cutoff frequency $f_{c}$ and order $o_{\text{LPF}}$. The cutoff frequency controls the width of the passband, which affects the range of frequencies with the highest power density in the drawn control signal perturbations. A lower $f_{c}$ results in a greater bias at lower frequencies and an increased use of low-frequency signals. In turn, a higher $f_{c}$ reduces the bias on low frequencies and allows searching in a wider range of signals. The second parameter of the Butterworth filter is its order $o_{\text{LPF}}$, which controls the slope of the power density spectrum in the stopband – roll-off, i.e., how much the attenuation of the signal components grows with the growth of the frequency. The higher the order, the lesser the exploration outside the passband, and the greater the bias towards low frequencies. To better visualize these dependencies, in Figure 3 we present the impact of the filter cutoff frequency $f_{c}$ and its order $o_{\text{LPF}}$ on the magnitude of the frequency response and on the signals in the time domain.
Last but not least, one of the benefits of the proposed approach is the interpretability of the cutoff frequency $f_{c}$ parameter, as it enables intuitive tuning. One of the natural approaches to choosing this parameter is to set it around the natural cutoff frequency of the system to be controlled with LP-MPPI, as inducing higher frequencies requires a significant effort to overcome the damping characteristics of the system itself. Moreover, one can often heuristically estimate the highest expected frequency of the system states necessary to maximize the reward or obtain the desired behavior. Furthermore, in some cases, the interpretability of the cutoff frequency may help regularize the system’s behavior, e.g., preventing exploitation of the carelessly designed reward function.
@@
To sum up, the proposed LP-MPPI approach filters out the high-frequency components from the evaluated control signals within the MPPI loop, significantly reducing the high-frequency components in the obtained controls. This reduction is particularly important for robotic systems, as it decreases their wear and biases the search for high-performance control signals towards a more promising area (consider the plausibility of finding a performant control signal in the upper left and bottom right corners of the bottom of Figure 3). Finally, our approach enables one to effectively control the frequency spectrum of the control signal perturbations, while providing intuitive tuning.
@@ IV-A Gymnasium environments
Despite higher rewards, an important aspect of the control algorithm is the smoothness of the applied controls. To visualize the differences between the considered methods, we present the applied control trajectories for the Ant-v5 environment with $H=15$ and $N=100$ in Figure 6. For fairness of comparison, each algorithm was used with the optimal set of parameters, w.r.t. cumulative reward, determined with Optuna [1] using 100 trials. One can see that the original MPPI algorithm generates undesirably sharp controls, while the remaining baselines produce significantly smoother controls, comparable to each other and to our proposed LP-MPPI. To quantitatively assess smoothness, we report Mean Squared Second Derivative (MSSD) and Mean Savitzky-Golay Filter Deviation (MSGFD) in Table I. The results indicate that our method attains very low MSSD and the smallest MSGFD, confirming that the control trajectories it generates are notably smooth.
Figure 6: Sample applied control signals in the Ant-v5 environment for the considered MPPI-based control algorithms.
@@ IV-B Quadruped locomotion
In the previous experiment, we showed that LP-MPPI outperforms the baselines for simplified abstract robots. Instead, in this one, we focus on the full-scale simulated quadrupeds with 12DoF. We consider a locomotion task with trot gait, imposed in the reward function to guide the search of the control signals, on two quadrupeds, i.e., Unitree Go2 and MAB Silver Badger (see Figure 4).
Recently, it was shown that MPPI-based approaches can succeed in such complex control tasks by diffusion-style annealing of the standard deviation of the noise distribution and the use of spline interpolation, as it was done in the Dial-MPC approach [28]. We would like to enhance the Dial-MPC with the proposed low-pass filtering (LP-Dial-MPC) to evaluate whether it applies to recent MPPI-based approaches and can improve their performance. Moreover, we would like to compare our method, in this setting, with the best-performing approach from the previous experiment – ColoredMPPI [24]. Thus, we extend the Dial-MPC with colored noise instead of the default white one.
@@
In this experiment, to highlight the robustness of the proposed approach to the choice of its parameters, we do not perform the search with Optuna but instead report the performance for several intuitive parameter sets, i.e. cutoff frequency $f_{c}\in\{2,3,4\}$ for Unitree and $f_{c}\in\{3,4,5\}$ for MAB robot, and orders $o_{\text{LPF}}\in\{2,3,4\}$. We compared their performance with the two above-mentioned baselines, for which we found the sets of the best parameters using Optuna. The results of this experiment can be found in Figure 7. The proposed low-pass filtering approach implemented into the Dial-MPC framework consistently outperforms the default Dial-MPC, by 24% and 41% for the Unitree Go2 and MAB Silver Badger robots, respectively. We attribute these improvements to the more fine-grained control (higher number of decision variables) of the LP Dial-MPC and its ability to directly shape the frequency spectrum of the sampling distribution. In turn, the spectrum shaping capabilities of the colored noise are relatively limited and bias only the lowest frequencies, which results in significantly worse performance (about two times lower rewards than ours).
@@ IV-C Real-world F1TENTH racing
In all previous experiments, we assumed that the models of the controlled systems are perfectly known and are used by the MPPI to search for the best control trajectories. In turn, in this task, we would like to use an analytical model of the F1TENTH car (dynamic single-track model [26] with MF6.1 tire model [7]) and evaluate it under the real-world racing conditions (see Figure 4). The goal of this task is to cover the highest possible distance around the track centerline in 30 s, on the $14.2\text{\,}\mathrm{m}$ long $1\text{\,}\mathrm{m}$ wide oval racetrack. We defined the cost function by
@@
where $v_{f}$ is the velocity along the centerline, $n$ is the distance to the centerline, $T_{w}$ is the track width, $\alpha$ is the slip angle, $\theta$ is the vehicle orientation, and $T_{\theta}$ is the orientation of the centerline. We set the horizon $H=30$, $dt=50$ ms, control frequency to $30\text{\,}\mathrm{Hz}$, and evaluated the algorithms for both $N=10$ and $N=50$ rollouts. We have chosen the number of rollouts and the horizon to meet real-time requirements with a single core of Intel Core i5-12500H CPU, while achieving reasonable driving performance (see video attachment). The parameters of all methods were chosen in simulation using Optuna [1] with 50 trials.
@@
Figure 8: Distances [m] covered by the F1TENTH car in 30-second time trials using different MPPI algorithm variants. The proposed LP-MPPI achieves the highest median distances both for $N=10$ and $N=50$ rollouts.
@@ IV-D Computational overhead
An important aspect of every control method is its computational efficiency. Therefore, we evaluate how much computational overhead the proposed method introduces relative to the nominal MPPI and how it relates to the other considered baselines. To do so, we run all methods for 10 episodes in Ant-v5 (MuJoCo model, $H=15,N=100$) and F1TENTH (compiled analytical model, $H=30,N=10$) environments, using a single core of the Intel Core i5-12500H CPU. We compute the median of the control computation time and relate it to the one obtained by the MPPI. In Table II, we present the results of this experiment. One can see that all baselines introduce some notable computational overhead in the F1TENTH environment, since the compiled analytical model, which is responsible for most of the computations, is very fast. Note that the proposed method introduces the second smallest overhead of 2.4%. In turn, in the case of a relatively heavy dynamics model, e.g., Ant-v5 environment, we observe some counterintuitive results, like the decrease in the compute time for LP-MPPI and ColoredMPPI. We suppose that this may be caused by the variability in the timings, due to the use of a standard OS instead of a real-time one, or be an effect of filtering out the higher frequencies from the control signal, which may simplify the underlying physics simulation. In summary, a computationally intense dynamics evaluation causes the overhead introduced by the proposed method to be negligible.
@@ V Conclusions
In this work, we introduced Low-Pass Model Predictive Path Integral Control (LP-MPPI), a novel and easy-to-implement enhancement to MPPI that incorporates low-pass filtering into the sampling process. By directly shaping the frequency spectrum of control trajectory perturbations, LP-MPPI eliminates harmful high-frequency noise and improves the efficiency of searching for optimal control trajectories. Unlike existing smoothing techniques or colored noise sampling, our approach offers intuitive fine-grained control over the optimal control search in the frequency domain, making it highly adaptable to various robotic systems.